# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v2

@@ 3.3. Insight
1. Explicit prefix identification and sharing. Instead of hashing and matching the common prefix at runtime and manage the prefix with LRU cache, BatchLLM explicitly identifies the common prefix globally within the large batch ahead-of-time, which will not miss the opportunity of prefix sharing caused by the implicit cache. Besides, BatchLLM refactors the prefix tree with a Dynamic Programming algorithm to enlarge the first-level prefix to avoid the system complexity and kernel overhead of the multi-level prefix.
@@
2. Grouped scheduling, request reordering and memory-centric token batching. BatchLLM schedules the requests at the granularity of a group of requests sharing the common prefix, which makes the prefix sharing convenient and shrinks the lifetime of the prefix’s KV memory. BatchLLM reorders the requests according to the length of the prompt and the estimated decoding length. As indicated in Fig.1, BatchLLM will schedule the requests with larger ratio of decoding length to prompt length with higher priority. In this way, the longer prompts will be scheduled later and can be better mixed with the earlier decoding tokens. It also forms the token-batches with the consideration of the KV memory usage, allowing to batch more tokens together to increase the size of token-batches and reduce the "valleys" in Fig.2.
3. Attention kernel optimization with horizontal fusion. BatchLLM horizontally fuses the computation on different KV chunks into the same kernel to reduce the tail effect and kernel launch overhead.
4. Design Methodology
4.1. Overview
Figure 3. BatchLLM overview.
@@ 4.1. Overview
Fig.3 shows the overview of BatchLLM optimizations. First, it analyzes the large batch of input prompts and identifies the common prefix explicitly. This is done before the scheduling of the requests. It simplifies the multi-level prefix into a single level prefix with the insight that the prefixes of the industry tasks are usually dominated by a long context (e.g., web document), as discussed in Sec.3.1. The explicit prefix processing is illustrated in Sec.4.2. The requests will be organized into groups where each group corresponds to the queries sharing the same prefix. The group will be the basic unit of task scheduling. Before scheduling the groups, BatchLLM will also reorder the groups according to the ratio of prefill length and postpone the groups with longer prefill. It then forms the token-batches with the consideration of the KV memory usage. This aims to better mix the decoding steps with the prefill chunks to increase the overall token-batch size. The throughput-oriented scheduling and token-batching optimization is described in Sec.4.3. BatchLLM also implements the horizontal fused Attention kernel to reduce the tail effect and kernel launch overhead, described in Sec.5.
4.2. Explicit Global KV Reuse Identification
@@
As discussed in Sec.3.1, different prompts may have multiple levels of prefixes. The multi-level prefixes can lead to system challenges and overhead of token batching and fused Attention kernel. For the token batching, it requires to manage the dependencies of the different levels of prefixes. For the Attention kernel, it will introduce more GPU kernels of the multi-level reuse. Instead, BatchLLM compresses the multi-level prefixes into a single level prefix to avoid the above challenges. This is based on the insight that the length of the prefixes are usually dominated by a single level due to the stronger reasoning ability of the newer LLMs and the task-specific fine-tuning (e.g., the prefix corresponding to the web document body can be much longer than other levels of the prefix, as discussed in Sec.3.1).
@@
As discussed in Sec.3.3, BatchLLM explicitly identifies the common prefix among the prompts within the large batch. It makes use of the compact prefix tree [^5] The compact prefix tree, also called Radix tree, is the prefix tree in which each node that is the only child is merged with its parent. to identify and represent the common prefix between prompts. However, directly using the first level prefix of the basic prefix tree as the common prefix can lead to suboptimal prefix reusing. Take the pesudo prompts in Fig.4d as an example. The 6 prompts are organized with the compact prefix tree in Fig.4a, where each node represents several consecutive tokens of the prompt and the tokens in a node is shared by its children. The right child of the root node (i.e., token ’b’) is shared by 5 prompts in total (prompt 2 to 6) according to Fig.4a, thus reusing the first level prefix can save 4 tokens’ KV computation. However, it is easy to infer that sharing the first 8 tokens between prompt 4 and 5 can save 8 tokens’ KV computation. Thus it requires to refactor the naive prefix tree to maximize the achieved first-level token reusing.
@@
1$\triangleright$ The 1st-level prefixes of a node are its children’s tokens
@@
6:         MaximizeReuse($child$) $\triangleright$ Solve sub-problems
@@
8:              $gain$leftarrow (leaves($gchild$) - 1) $\times$ tokens($gchild$)
9:              $penalty$leftarrow tokens($child$)
10:              if $gain$ > $penalty$ then
11:                  Fork $child$ and merge with $gchild$
@@
We design a Dynamic Programming algorithm on the compact prefix tree to maximize first-level token reusing, shown in Algo.1. The basic insight is that, given a node $D$ to maximize its first-level token reusing, it first solves the sub-problems of all $D$’s children to maximize each child’s first-level token reusing (line 6 in Algo.1), it then try to merge $D$’s first-level prefix with its children’s to see if it can enlarge $D$’s token reusing (line 7-11 in Algo.1). It recursive performs the above procedure until reaching the root note, by calling MaximizeReuse(root). Fig.4b and Fig.4c shows an example of this procedure. In Fig.4b, the left node in level 3 of the tree is forked and merged with it’s right child, which enlarges the 1st-level prefix reusing of the right node in level 2. Similarly, the right node in level 2 is forked and merged with its second child to enlarge the 1st-level prefix reusing of the root node. After maximizing the first-level prefix, the other levels of prefixes will be expanded during the token batching and will not be regarded as the shared context.
@@
We have compared the token saving ratio between the original multi-level prefix and the enlarged single prefix. The saving ratio is calculated according to Eq.2. For a dataset of the snippet generation task, the saving ratio of the original multi-level prefixes is about 56%, and the ratio is about 55% for the first-level prefix after enlarging (Sec.6.2.2). It means the enlarged first-level prefix only has 1% loss of the token reuse compared to the multi-level prefix for this typical workload, with the advantage of reducing the complexity and the overhead of the token batching and Attention kernel optimization. Besides, the DP algorithm only introduces less than 0.01% overhead in our evaluation (Sec.6.3.4).
Figure 4. The preprocessing to maximize the first level prefix reusing. It converts from the initial prefix tree (a) to the final prefix tree (c). Each cycle represents a node containing a number of tokens. The number in the cycle is the token number of the node. It iterates bottom-up to maximize the first level reusing recursively until reaching the root node.
4.3. Throughput-oriented Token Batching
4.3.1. Prefix-sharing Group Based Scheduling
@@
As described in Sec.4.2, BatchLLM identifies the common prefix explicitly ahead-of-time. To simplify the KV reuse and reduce the lifetime of the common prefix in the memory, BatchLLM schedules the requests sharing the same prefix all together. Specifically, the procedure described in Sec.4.2 will generate the prefix-sharing groups and BatchLLM will schedule the requests at the granularity of prefix-sharing group rather than each single request. A prefix-sharing group corresponds to the collection of requests sharing the same prefix. In this way, the requests sharing the same prefix can start at the same time and the memory of the common prefix can be released as soon as the group is completed. This differs from the LRU cache based prefix management (Kwon et al., 2023; Zheng et al., 2023) and can make sure all the common prefix can be reused the best and the memory will not be retained unnecessarily.
@@
To realize the prefix-sharing group based scheduling, BatchLLM maintains three queues of the to-be-batched tokens: common prefix queue, distinct prompt queue, and decoding queue. This method differs from the existing methods that use the same queue for all prefill tokens. At each iteration, BatchLLM will form the token-batch by fetching the tokens from the three queues in the order of decoding queue, distinct prompt queue, and common prefix queue (one constraint is that the a common prefix should be fetched and processed before its corresponding distinct prompts). On the one hand, fetching the decoding tokens before prefills can help to better mix the two types of tokens together in the token-batch (Holmes et al., 2024). On the other hand, fetching the decoding and distinct prompt tokens before the prefix tokens can make sure the active requests (i.e., the request that has been partially processed, either partial prefill chunks has been processed or already in decoding phase) can be scheduled before the inactive requests, which helps to finish the active requests earlier and thus release their memory earlier.
4.3.2. Request Groups Reordering
@@
As discussed in Sec.3.3, BatchLLM reorders the input requests to schedule the requests according to the ratio of prompt length to decoding length ($R$). We define the ratio $R$ as:
(3) $R=\frac{L_{decode}}{L_{prefill}}$
@@
In Eq.3, $L_{prefill}$ is the number of prefill tokens and $L_{decode}$ is the length of the output tokens. The requests with larger $R$ will be scheduled earlier. This helps to issue the requests with long decoding steps earlier and make the latter prompts’ length larger and thus can be better mixed with the previous decoding tokens to enlarge the token-batch size (example in Fig.1). The challenge is that the exact number of output tokens is not known before finishing the execution. We profiled the dataset of several typical information processing tasks (snippet generation, offline ranking, ads, etc.) and observed that the distribution of output length is relatively concentrated for a task. We thus assume the output length is constant and normalize $L_{prefill}$ into 1 in Eq.3 for all requests uniformly. As we schedule the requests in the unit of prefix-sharing group, the ratio of the group (after normalizing the length of output) becomes:
(4) $R_{group}=\frac{1}{L_{prefix}+\sum L_{distinct}}$
@@
According to Eq.4, the request groups with larger $R_{group}$ will be scheduled with higher priority.
4.3.3. Memory-centric Token Batching to Saturate GPU
@@
The request group reordering described above helps to reduce the "valleys" in the timeline of the token numbers ("valley" example in Fig.2) by providing more chances for the decoding tokens to be able to find the prefill chunks to mix together. Another limiter of the size of token-batches is that, the existing works (Kwon et al., 2023) use a fixed request number to limit the token batching, i.e., the request number cannot exceed a fixed threshold within a token-batch. For a token-batch dominated by the decoding tokens, it is easy to reach the upper bound of the request number while still have not achieved a large number of the total tokens. For example, vLLM sets the threshold of request number as 256 by default, if a token-batch already has 256 decoding tokens, it will have no chance to add more prefill chunks into this token-batch even if there is still enough memory to hold the KV memory of the prefill chunks. Note that purely increasing this threshold does not solve the problem in practice, as noted by the vLLM community666 https://github.com/vllm-project/vllm/issues/6801. This is because directly excessively enlarging batch size may cause the engine to have more frequent memory swaps and recomputation due to the GPU memory can be not enough to hold the new coming KV tensor.
@@
5. Implementation
Horizontal Fusion of Prefix-shared Attention. As described in Sec.3.2, the existing works use separate kernels to compute the partial Attention on the prefix and the distinct KV context. This will lead to the tail effect of each kernel and the launch overhead of these kernels. BatchLLM horizontally fuses the two partial Attention calculation into one kernel. Different parts of the computation are performed in different thread blocks. Note that the problem shape of the two parts are different. BatchLLM applies different tiling configuration for the two parts to achieve the best performance. It uses the common auto-tuning method to find the best configuration. We implement the Attention in BatchLLM through OpenAI Triton language (Tillet et al., 2019). The limitation of the Triton based implementation is that, the performance of Triton version of FlashAttention is still worse than that of CUDA version (Sec.6.3.3). However, the advantage is that Triton supports multiple platforms (both NVIDIA and AMD GPUs) so that BatchLLM implementation can run on both NVIDIA and AMD GPUs. We use the builtin autotuner of Triton to tune the tiling size of the Attention implementation in BatchLLM. Besides that, some ahead-of-time compilation methods are applied to reduce the launching overhead, including warm-up process on NVIDIA GPUs, and AOTriton [^7] https://github.com/ROCm/aotriton on AMD GPUs.
We implement the prefix identification (Sec.4.2) in Python as a standalone preprocessing script. It accepts a list of inputs and generate a list of prefix-sharing groups. The inputs can be the nested input ids (e.g., [<list of input 0 tokens>, <list of input 1 tokens>, ...]), and each prefix-sharing group is represented as a tuple of the prefix and the corresponding distinct prompts (e.g., tuple(<list of prefix tokens>, [<list of distinct 0 tokens>, <list of distinct 1 tokens>, ...])).
We integrate the token batching optimization (Sec.4.3) in vLLM (Kwon et al., 2023). Specifically, we customize the vLLM to accept the list of prefix-sharing group tuples generated by the preprocessing script, and implement the group-wised scheduling and token batching logic upon the vLLM token batching function.
(a) Llama 3.1 8B throughput on A100, request-count = 6400
(b) Qwen 2.5 14B throughput on A100, request-count = 6400
(c) Llama 3.1 70B throughput, on A100, request-count = 800
(d) Llama 3.1 8B throughput, on MI200, request-count = 6400
(e) Qwen 2.5 14B throughput, on MI200, request-count = 6400

### v3

@@ 3.1 Overview
1. Explicit prefix identification and sharing. Instead of managing the prefix with LRU cache, BatchLLM explicitly identifies the common prefix globally within the large batch ahead-of-time, which will not miss the opportunity of prefix sharing caused by the implicit cache. Besides, BatchLLM refactors the prefix tree with a Dynamic Programming algorithm to enlarge the first-level prefix to avoid the system complexity and kernel overhead of the multi-level prefix.
@@
2. Grouped scheduling, request reordering and memory-centric token batching. BatchLLM schedules the requests at the granularity of a group of requests sharing the common prefix, which makes the prefix sharing convenient and shrinks the lifetime of the prefix’s KV memory. BatchLLM reorders the requests according to the length of the prompt and the estimated decoding length. As indicated in Fig.1, BatchLLM will schedule the requests with larger ratio of decoding length to prompt length with higher priority. In this way, the longer prompts will be scheduled later and can be better mixed with the earlier decoding tokens. It also forms the token-batches with the consideration of the KV memory usage, allowing to batch more tokens together to increase the size of token-batches and reduce the "valleys".
Figure 3: BatchLLM overview.
@@
Fig.3 shows the overview of BatchLLM optimizations. First, it analyzes the large batch of input prompts and identifies the common prefix explicitly. This is done before the scheduling of the requests. It simplifies the multi-level prefix into a single level prefix with the insight that the prefixes of the industry tasks are usually dominated by a long context (Sec.2.1). The explicit prefix processing is illustrated in Sec.3.2. The requests will be organized into groups where each group corresponds to the queries sharing the same prefix. The group will be the basic unit of task scheduling. Before scheduling the groups, BatchLLM will also reorder the groups according to the ratio of prefill length and postpone the groups with longer prefill. It then forms the token-batches with the consideration of the KV memory usage. This aims to better mix the decoding steps with the prefill chunks to increase the overall token-batch size. The throughput-oriented scheduling and token-batching optimization is described in Sec.3.3.
3.2 Explicit Global KV Reuse Identification
@@
As discussed in Sec.2.1, different prompts may have multiple levels of prefixes. The multi-level prefixes can lead to system challenges and overhead of token batching and fused Attention kernel. For the token batching, it requires to manage the dependencies of the different levels of prefixes. For the Attention kernel, it will introduce more GPU kernels of the multi-level reuse. Instead, BatchLLM compresses the multi-level prefixes into a single level prefix to avoid the above challenges, based on the insight that the length of the prefixes are usually dominated by a single level (Sec.2.1).
Figure 4: The preprocessing to maximize the first level prefix reusing. Each circle represents a node containing a number of tokens. The number in the circle is the token number of the node.
@@
BatchLLM explicitly identifies the common prefix among the prompts within the large batch. It makes use of the compact prefix tree [^2] The compact prefix tree, also called Radix tree, is the prefix tree that merges the only child with its parent. to identify and represent the common prefix between prompts. However, directly using the first level prefix of the basic prefix tree as the common prefix can lead to suboptimal prefix reusing. Take the pseudo prompts in Fig.4d as an example. The 6 prompts are organized with the compact prefix tree in Fig.4a, where each node represents several consecutive tokens of the prompt and the tokens in a node is shared by its children. The right child of the root node (i.e., token ’b’) is shared by 5 prompts in total (prompt 2 to 6) according to Fig.4a, thus reusing the first level prefix can save 4 tokens’ KV computation. However, it is easy to infer that sharing the first 8 tokens between prompt 4 and 5 can save 8 tokens’ KV computation. Thus it requires to refactor the naive prefix tree to maximize the achieved first-level token reusing.
@@
[…]
@@
6:   MaximizeReuse(childchild) $\triangleright$ Solve sub-problems
@@
8:    gaingain $\leftarrow$ (leaves(gchildgchild) - 1) $\times$ tokens(gchildgchild)
9:    penaltypenalty $\leftarrow$ tokens(childchild)
10:    if gaingain > penaltypenalty then
11:     Fork childchild and merge with gchildgchild
@@
We design a Dynamic Programming algorithm on the compact prefix tree to maximize first-level token reusing, shown in Algo.1. The basic insight is that, given a node $D$ to maximize its first-level token reusing, it first solves the sub-problems of all $D$’s children to maximize each child’s first-level token reusing (line 6 in Algo.1), it then try to merge $D$’s first-level prefix with its children’s to see if it can enlarge $D$’s token reusing (line 7-11 in Algo.1). It recursive performs the above procedure until reaching the root note, by calling MaximizeReuse(root). Fig.4b and Fig.4c shows an example of this procedure. In Fig.4b, the left node in level 3 of the tree is forked and merged with it’s right child, which enlarges the 1st-level prefix reusing of the right node in level 2. Similarly, the right node in level 2 is forked and merged with its second child to enlarge the 1st-level prefix reusing of the root node. After maximizing the first-level prefix, the other levels of prefixes will be expanded during the token batching and will not be the shared context.
@@
We have compared the token saving ratio between the original multi-level prefix and the enlarged single prefix. The saving ratio is calculated according to Eq.1. For a dataset of the snippet generation task, the saving ratio of the original multi-level prefixes is about 56%, and the ratio is about 55% for the first-level prefix after enlarging (Sec.4.2.2). It means the enlarged first-level prefix only has 1% loss of the token reuse compared to the multi-level prefix for this typical workload, with the advantage of reducing the complexity and the overhead of the token batching and Attention kernel optimization. More details will be presented in Sec.4.3.1.
Time and Space Complexity Analysis. The complexity of Algo.1 is $O(nc^{2})$, where $n$ is the number of nodes and $c$ is the number of average children. The algorithm visits every node once and has $O(n)$ for traversal, and iterates over its children and their children for each node and has $O(c^{2})$ per node as the worst case. Thus, the time complexity is $O(nc^{2})$. In real scenarios, $c$ is typically small on average, so this is close to $O(n)$ in practice. The space complexity is $O(n)$. The memory overhead introduced by "fork child and merge with grandchild" is very small, as the metadata of the node is the token_id, which is in int64 datatype. The worst case is that all nodes are fully forked, which means the same prefix of different prompts are stored independently. In this case, the storage will be equal to the sum of token numbers of all prompts (in int64 datatype), which is still small compared to the memory consumption of model weights. Algo.1 does not introduce KV residency time as it is computed based on the input token_ids before processing the tokens. As a result, the DP algorithm only introduces less than 0.01% overhead in our evaluation (Sec.4.3.1)
BatchLLM does not completely disable the LRU mechanism. It explicitly applies global prefix cache to the longest first-level prefix and relies on vLLM’s LRU to manage the possible remaining prefixes (e.g., the second-level sub-prefix) for distinct prompt parts.
3.3 Throughput-oriented Token Batching
3.3.1 Prefix-sharing Group Based Scheduling
@@
To simplify the KV reuse and reduce the lifetime of the common prefix in the memory, BatchLLM schedules the requests sharing the same prefix all together. Specifically, the procedure described in Sec.3.2 will generate the prefix-sharing groups (the collection of requests sharing the same prefix) and BatchLLM will schedule the requests at the granularity of prefix-sharing group rather than each single request. In this way, the requests sharing the same prefix can start at the same time and the memory of the common prefix can be released as soon as the group is completed. This differs from the LRU cache based prefix management and can make sure all the common prefix can be reused the best and the memory will not be retained unnecessarily.
@@
To realize the prefix-sharing group based scheduling, BatchLLM maintains three queues of the to-be-batched tokens: common prefix queue, distinct prompt queue, and decoding queue. At each iteration, BatchLLM will form the token-batch by fetching the tokens from the three queues in the order of decoding queue, distinct prompt queue, and common prefix queue (one constraint is that the a common prefix should be fetched and processed before its corresponding distinct prompts). On the one hand, fetching the decoding tokens before prefills can help to better mix the two types of tokens together in the token-batch Holmes et al. (2024). On the other hand, fetching the decoding and distinct prompt tokens before the prefix tokens can make sure the active requests (i.e., the request that has been partially processed) are scheduled before the inactive requests, which finishes the active requests earlier and releases their memory earlier.
3.3.2 Request Groups Reordering
@@
BatchLLM reorders the input requests to schedule the requests according to the ratio of prompt length to decoding length ($R$). We define the ratio $R$ as:
$R=\frac{L_{decode}}{L_{prefill}}$ (2)
@@
In Eq.2, $L_{prefill}$ is the number of prefill tokens and $L_{decode}$ is the length of the output tokens. The requests with larger $R$ will be scheduled earlier. This helps to issue the requests with long decoding steps earlier and make the latter prompts’ length larger and thus can be better mixed with the previous decoding tokens to enlarge the token-batch size (example in Fig.1). The challenge is that the exact number of output tokens is not known before finishing the execution. We profiled the dataset of several typical industry tasks (ads, caption generation, offline ranking, etc.) and observed that the distribution of output length is relatively concentrated for the same kind of task, which means the effect of the output length in Eq.2 is less significant for the same kind of task as it does not vary a lot. We thus make an approximation by normalizing $L_{decode}$ into 1 in Eq.2 for all requests uniformly. As we schedule the requests in the unit of prefix-sharing group, the ratio of the group (after normalization) becomes:
$R_{group}=\frac{1}{L_{prefix}+\sum L_{distinct}}$ (3)
@@
The request groups with larger $R_{group}$ will be scheduled with higher priority.
(a) Output length distribution of Ads title rewrite.
(b) Output length distribution of caption generation.
Figure 5: Generation output length distribution of two typical real-industry workloads.
Fig.5 shows the output length distributions of two representative industry workloads. It shows that the the output lengths are in a very narrow range, which motivates our choice of normalizing decoding length.
3.3.3 Memory-centric Token Batching to Saturate GPU
@@
Another limiter of the size of token-batches is that, the existing works Kwon et al. (2023) use a fixed request number to limit the token batching, i.e., the request number cannot exceed a fixed threshold within a token-batch. For a token-batch dominated by the decoding tokens, it is easy to reach the upper bound of the request number while still have not achieved a large number of the total tokens. For example, vLLM sets the threshold of request number as 256 by default, if a token-batch already has 256 decoding tokens, it will have no chance to add more prefill chunks into this token-batch even if there is still enough memory to hold the KV memory of the prefill chunks. Note that purely increasing this threshold does not solve the problem in practice (noted by the vLLM community [^3] https://github.com/vllm-project/vllm/issues/6801), as directly excessively enlarging batch size may cause the engine to have more frequent memory swaps and recomputation due to the GPU memory can be not enough.
@@
3.4 End-to-end System Integration
The integration of BatchLLM into other frameworks is quite easy. BatchLLM comprises two pluggable modules: (i) a token-batching scheduler and (ii) an attention backend, that can be added alongside existing components in common inference frameworks (e.g., vLLM, SGLang) without altering user-facing APIs or the serving loop. Token-batching scheduler is orthogonal to the framework’s existing continuous batching/prefill–decode pipeline. We add lightweight per-request metadata to indicate shared prefix vs. distinct part. The scheduler ensures the common prefix across requests is computed once and reused, reducing redundant calculation in both attention (KV-cache construction) and non-attention compute (e.g., projection and FFN GEMMs) during prefill and decoding. Attention backend is a drop-in backend with full fallback. The backend presents the same public interface as the framework’s attention module and simply registers as an additional backend. If shapes/hardware are unsupported, it falls back to the standard kernel, ensuring compatibility and graceful degradation.
(a) Llama 3.1 8B throughput on A100, #requests = 6400
(b) Qwen 2.5 14B throughput on A100, #requests = 6400
(c) Llama 3.1 70B throughput, on A100, #requests = 800
(d) Llama 3.1 70B throughput, on A100-TP2, #requests = 800
(e) Llama 3.1 8B throughput with Poisson perturbed lengths on A100, #requests = 6400
(f) Llama 3.1 8B throughput, on MI200, #requests = 6400
(g) Qwen 2.5 14B throughput, on MI200, #requests = 6400